# Learning Neural Control Barrier Functions from Offline Data with Conservatism

### v1

@@ Abstract
Safety filters, particularly those based on control barrier functions, have gained increased interest as effective tools for safe control of dynamical systems. Existing correct-by-construction synthesis algorithms, however, suffer from the curse of dimensionality. Deep learning approaches have been proposed in recent years to address this challenge. In this paper, we contribute to this line of work by proposing an algorithm for training control barrier functions from offline datasets. Our algorithm trains the filter to not only prevent the system from reaching unsafe states but also out-of-distribution ones, at which the filter would be unreliable. It is inspired by Conservative Q-learning, an offline reinforcement learning algorithm. We call its outputs Conservative Control Barrier Functions (CCBFs). Our empirical results demonstrate that CCBFs outperform existing methods in maintaining safety and out-of-distribution avoidance while minimally affecting task performance.
Keywords: Safety filters, Control barrier functions, Offline reinforcement learning
Figure 1: We propose a simple modification of the training process of neural control barrier functions when used for learning from offline data. Our learned functions can be used to define constraints in quadratic programs that can be used as safety filters that correct unsafe and out-of-distribution nominal controls. We evaluate our approach in four environments and demonstrate its improved performance over three baseline methods.
1 Introduction
@@
Ensuring that autonomous systems satisfy safety specifications is critical to their deployment in various settings, such as autonomous driving [1] and robotic surgery [2]. Safety filters have emerged as promising tools for enforcing safety constraints [3]. Control barrier functions (CBF) allow the design of efficient safety filters in the form of quadratic programs (QP) that preserve the invariance of safe sets while minimally altering nominal control [4, 3, 5]. However, synthesizing CBFs using correct-by-construction techniques, such as sum-of-squares programming [6, 7], do not scale beyond few dimensions.
@@
To address this challenge, researchers have proposed data-driven methods for learning safety filters [8, 9, 10], which have been applied in settings with uncertain dynamics [11, 12] and high-dimensional observations, such as images [10, 13, 14, 15, 16, 17, 18] and point clouds [19, 20]. However, these filters lack the theoretical guarantees that their formally verified or synthesized counterparts have. Moreover, when deployed, by design, they will alter the nominal controls resulting in a shift between the distribution of trajectories used for their training and the ones visited during deployment. It is often assumed that the filter is trained iteratively until convergence: in each iteration, it is used to generate new trajectories whose states get labeled as safe or unsafe, and then the filter gets retrained accordingly [21]. After convergence, the distribution of states used to train the filter would match the distribution of states it will encounter during deployment. Consequently, one can use classic generalization bounds to obtain probabilistic guarantees on the correctness of the learned filter. However, in various safety-critical settings, it can be challenging and expensive to collect new trajectories, especially since they may be unsafe when generated using the filter being trained. This highlights the need for learning safety filters from offline data, while addressing or mitigating the resulting distribution shift between training and deployment.
@@
Offline reinforcement learning (RL) has been studied extensively in the past few years [22]. Several algorithms have been proposed to learn optimal policies from offline datasets of trajectories in that literature [23, 24, 25, 26, 27, 28]. The goal is to obtain policies that outperform the ones used to generate the data or those which can be obtained using behavior cloning (BC), while accounting for distribution shift. In this paper, we aim to train safety filters using offline datasets of safe and unsafe trajectories that alter unsafe nominal controls to in-distribution (InD) safe ones without being overly restrictive. For that purpose, we introduce Conservative Control Barrier Functions (CCBFs), a method for training neural CBFs that is inspired by conservative Q-learning (CQL) [23], a prominent offline RL algorithm. The core insight is that the learned CBF should assign higher safety scores to InD states than to out-of-distribution (OOD) states, even if both are safe. This objective is similar to the CQL objective, which penalizes the overestimation of Q-values for OOD actions. By analogously discouraging high safety scores for OOD states, safety filters using CCBFs keep the system in the region of the state space in which they can provide confident estimates of safety.
@@
2 Related Work
@@
Several algorithms have been proposed recently to train neural CBFs from offline datasets [29, 30, 31, 12, 32, 8, 33, 34, 35, 36]. They differ in their objectives, their dataset generation process, their assumptions, and their guarantees, if any. Castaneda et al. [37] and Kang et al. [38] proposed learning safety filters to prevent control systems from reaching OOD states. Castaneda et al. [37] proposed an algorithm for training neural CBFs, termed in-distribution barrier functions (iDBFs), using an offline dataset of safe trajectories to prevent the control system from reaching OOD states. To train an iDBF, the algorithm first trains a BC policy using the same dataset. From each state in the training dataset, the algorithm samples random actions. If the BC policy assigns a sampled action a probability density below a user‑defined threshold $p$, the next state reached by that action is labeled OOD and appended to the dataset. It finally uses an existing approach for training neural CBFs (e.g., [11]) to train the iDBF using the augmented dataset, treating the InD states as safe and the OOD states as unsafe. Kang et al. [38] proposed Lyapunov Density Models, combining control Lyapunov functions [39] and density models [40], which are also learned from data to keep a control system from reaching OOD states. More recently, Yu et al. [41] proposed an algorithm for learning neural CBFs from an offline dataset consisting of a mix of labeled and unlabeled trajectories by leveraging cost-sensitive classification [42] to assign safety labels to the unlabeled demonstrations.
Several other works proposed algorithms for synthesizing CBFs from expert safe demonstrations providing formal correctness guarantees through Lipschitz continuity and coverage-based proofs [8, 43, 12]. These works assume that the dataset is sufficiently large and spread to cover the region of the state space that the system will operate in during deployment, and thus no generalization capabilities are needed from the learned filter and OOD states are not reachable.
In this paper, we aim to train neural CBFs from offline datasets of both safe and unsafe demonstrations that can be used as safety filters for control systems to avoid both unsafe and OOD states while minimally affecting task progress.
@@
Safe offline RL is a branch of offline RL that aims to train policies that satisfy safety constraints while maximizing reward. It tackles a similar problem to ours: given a set of trajectories of a Constrained Markov Decision Process (CMDP), design an optimal and safe policy that avoids distribution shift [44, 45]. However, it considers soft safety constraints which require the expected cost of the trajectories to remain below a user-defined threshold, in contrast with the hard constraints of avoiding unsafe states considered in our case. Several approaches incorporated safety filters into the safe offline RL pipeline [46, 47, 48]. A close work to ours is the recently proposed method FISOR [46], which similarly employs a hard state-wise constraint formulation for safe offline RL. It first adapts Hamilton Jacobi reachability to learn a value function that defines the backward reachable set of a known constraint set from the offline dataset. It uses Implicit Q-learning [24] to prevent over-optimistic safety value estimation that can result from querying OOD actions in offline RL settings. It then trains a weighted BC policy that maximizes rewards within the feasible region in the state space, i.e., the complement of the backward reachable set. We instead take a different approach and assume that the states in the dataset are labeled as safe or unsafe, train a neural CBF with an additional loss term inspired from CQL [23] to avoid over-optimistic CBF values for OOD states, and test its performance as a safety filter for various nominal controllers that are learned separately from the same dataset using various offline RL algorithms.
3 Preliminaries
3.1 Control barrier functions
We assume that the dynamics of the systems we consider in this paper are control-affine.
Definition 1 (Control-affine systems).
@@ 3.1 Control barrier functions
Definition 2 (Control barrier functions [4]).
@@
A function $B:\mathcal{X}\to\mathbb{R}$ is a control barrier function for system (1) if it is continuously differentiable and satisfies:
$\exists u\in\mathcal{U}\text{ such that }\dot{B}(x,u)+\gamma(B(x))\geq 0,\quad\forall x\in\mathcal{D}\subseteq\mathcal{X},$ (2)
@@
where the super-level set $B_{\geq 0}:=\{x\ |\ B(x)\geq 0\}$ satisfies $B_{\geq 0}\subseteq\mathcal{D}$, and $\gamma:\mathbb{R}\to\mathbb{R}$ is a locally Lipschitz extended class-$\mathcal{K}_{\infty}$ function, meaning it is strictly increasing and satisfies $\gamma(0)=0$.
@@
A CBF $B$ for system (1) specifies the controls that guarantee the forward invariance of $B_{\geq 0}$. That means that all the trajectories that start in $B_{\geq 0}$ and follow control signals that satisfy (2) at all time instants, remain within it at all times, as stated in the following theorem. If the set of unsafe states is disjoint from $B_{\geq 0}$ and the system starts from states in $B_{\geq 0}$, then the system can be kept safe by following such controls.
Theorem 1 ([4]).
@@
Any Lipschitz continuous control policy $\pi:\mathcal{D}\to\mathcal{U}$ satisfying
$\forall x\in\mathcal{D},\ \pi(x)\in K_{\textit{cbf}}:=\{u\in\mathcal{U}\mid\nabla B(x)(f(x)+g(x)u)+\gamma(B(x))\geq 0\}$
makes the set $B_{\geq 0}$ forward invariant.
@@
Given a reference controller $\pi_{\text{ref}}:\mathcal{X}\to\mathcal{U}$ that does not necessarily satisfy safety constraints, a safety filter using the CBF $B$ can be employed to modify its unsafe control choices. Specifically, a quadratic program (QP) can be formulated to determine the closest safe control to $\pi_{\text{ref}}(x)$ [5], resulting in what is usually denoted by a CBF-QP policy $\pi_{\text{safe}}$, as shown below:
$\pi_{\text{safe}}(x)$ $:=\arg\min_{u\in\mathcal{U}}\lVert u-\pi_{\text{ref}}(x)\rVert^{2}\quad s.t.\quad\nabla B(x)(f(x)+g(x)u)+\gamma(B(x))\geq 0.$ (3)
3.2 Offline reinforcement learning
Offline reinforcement learning enables designing optimal control policies from offline datasets that avoid OOD states and actions. CQL [23] is an offline RL algorithm that penalizes the $Q$-values of state-action pairs that do not appear in the offline dataset to prevent over-estimating their values.
The standard Q-learning objective based on the Bellman equation is to minimize the following loss function: $\mathcal{L}_{base}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{\theta}(s,a)-\left(r+\gamma\max_{a^{\prime}}Q_{\theta}(s^{\prime},a^{\prime})\right)\right)^{2}\right]$, where $\mathcal{D}$ is the offline dataset, $Q_{\theta}$ is the $Q$ function being learned with parameters $\theta$, and $\gamma$ is the discount factor [49]. CQL augments this function with another term that aims to prevent over-estimating the $Q$-values of OOD data as follows: $\mathcal{L}_{CQL}(\theta)=\mathcal{L}_{base}(\theta)+\alpha\cdot\mathcal{L}_{cons}(\theta),$ where $\alpha$ is a hyperparameter and $\mathcal{L}_{cons}$ is a regularization term that penalizes the $Q$-values of OOD state-action pairs. In practice, $\mathcal{L}_{cons}$ can be implemented in various ways, one of which is the following: $\mathcal{L}_{\text{cons}}(\theta)=\mathbb{E}_{s\sim\mathcal{D}}\bigl{[}\log\sum_{a}\exp(Q_{\theta}(s,a))-\mathbb{E}_{a\sim\hat{\pi}_{\beta}(\,\cdot\mid s)}[Q_{\theta}(s,a)]\bigr{]}$, where $\hat{\pi}_{\beta}(a\mid s)$ is the empirical behavior policy, i.e., an estimate of the conditional distribution of actions given state $s$ induced by the offline dataset $\mathcal{D}$. The first part of the term, $\log\sum_{a}\exp(Q_{\theta}(s,a))$, is a soft maximum over the $Q$-values corresponding to all possible actions at a given state. Minimizing this term reduces the $Q$-values of all actions. The second component, $\mathbb{E}_{a\sim\hat{\pi}_{\beta}(\,\cdot\mid s)}[Q_{\theta}(s,a)]$, is the expected Q-value for actions followed at state $s$ in the trajectories in the dataset. Maximizing this term encourages assigning high $Q$-values for InD actions. As a result, unseen actions receive lower $Q$-values while in-distribution actions retain accurate estimates, so the derived policy $\pi(s)=\arg\max_{a}Q(s,a)$ would select in-distribution actions.
3.3 Offline safe reinforcement learning
Safe RL extends RL to settings where an agent has to keep the total cost below a predefined threshold while maximizing reward. It aims to train a policy $\pi$ for a Constrained Markov Decision Process (CMDP) that maximizes $\mathbb{E}_{s}[V^{\pi}_{r}(s)]$ subject to the constraints $\mathbb{E}_{s}[V^{\pi}_{c}(s)]\leq\kappa$ and $D(\pi\|\pi_{\beta})\leq\epsilon$, where $V^{\pi}_{r}(s)$ represents the reward value function, $V^{\pi}_{c}(s)$ represents the cost value function, $\kappa$ is a user-defined cost threshold, $\pi_{\beta}$ is the potentially-unknown behavioral policy that was used to generate the offline dataset $\mathcal{D}:=\{(s,a,s^{\prime},r,c)\}$, and $D(\pi\|\pi_{\beta})$ measures the divergence between $\pi_{B}$ and $\pi$.
3.4 Problem statement
We consider the setting where an offline dataset consisting of safe and unsafe trajectories is available. Our objective is to train a safety filter that prevents the system from reaching unsafe states. Consequently, it must also prevent the system from reaching OOD ones at which it might not be reliable enough to determines safe from unsafe states and actions. The safety filter should not be conservative beyond what is necessary to maintain safety, i.e., as minimally affecting nominal control as possible.
4 Method

### v2

# Learning Conservative Neural Control Barrier Functions from Offline Data

@@ Abstract
Safety filters, particularly those based on control barrier functions, have gained increased interest as effective tools for safe control of dynamical systems. Existing correct-by-construction synthesis algorithms for such filters, however, suffer from the curse-of-dimensionality. Deep learning approaches have been proposed in recent years to address this challenge. In this paper, we add to this set of approaches an algorithm for training neural control barrier functions from offline datasets. Such functions can be used to design constraints for quadratic programs that are then used as safety filters. Our algorithm trains these functions so that the system is not only prevented from reaching unsafe states, but is also disincentivized from reaching out-of-distribution ones, at which they would be less reliable. It is inspired by Conservative Q-learning, an offline reinforcement learning algorithm. We call its outputs Conservative Control Barrier Functions (CCBFs). Our empirical results demonstrate that CCBFs outperform existing methods in maintaining safety while minimally affecting task performance. Source code is available at https://github.com/tabz23/CCBF.
I Introduction
@@
Ensuring that control systems satisfy safety specifications is critical in various applications, such as autonomous driving [1] and robotic surgery [2]. Safety filters have emerged as promising tools for enforcing safety constraints [3]. Control barrier functions (CBF) allow the design of efficient safety filters in the form of quadratic programs (QP) that preserve the forward invariance of sets of states disjoint from a user-defined failure set of states, such as collisions, while minimally altering (potentially unsafe) task-achieving nominal control [4, 5]. However, synthesizing CBFs using correct-by-construction techniques, such as sum-of-squares programming [6, 7], do not scale beyond few dimensions.
@@
To address this challenge, researchers have proposed data-driven methods for learning safety filters [8], which have been applied in settings with uncertain dynamics [9, 10] and high-dimensional observations, such as images [11, 12, 13] and point clouds [14, 15]. However, these filters lack the theoretical guarantees that their formally verified or synthesized counterparts have. Moreover, when deployed, by design, they will alter the nominal controls resulting in a shift between the distribution of trajectories used for their training and the ones visited during deployment. It is often assumed that the filter is trained iteratively until convergence: in each iteration, it is used to generate new trajectories whose states get labeled as safe , i.e., belong to a forward invariant set disjoint from the failure set, or unsafe , i.e., belong to the complement of that forward invariant set, and then the filter gets retrained accordingly [16]. After convergence, the distribution of states used to train the filter would match the distribution of states it will encounter during deployment. Consequently, one can use classic generalization bounds to obtain probabilistic guarantees on the correctness of the learned filter. However, in various safety-critical settings, it can be challenging and expensive to collect new trajectories. This highlights the need for learning safety filters from offline data, while addressing or mitigating the resulting distribution shift between training and deployment.
@@
Offline reinforcement learning (RL) has been studied extensively in the past few years [17]. Several algorithms have been proposed to learn optimal policies from offline datasets of trajectories in that literature [18, 19, 20, 21, 22]. The goal is to obtain policies that outperform the ones used to generate the data or those which can be obtained using behavior cloning (BC), while accounting for distribution shift. In this paper, we aim to train safety filters using offline datasets of safe and unsafe trajectories that alter unsafe nominal controls to safe ones. For that purpose, we introduce Conservative Control Barrier Functions (CCBFs), a method for training neural CBFs that is inspired by conservative Q-learning (CQL) [18], a prominent offline RL algorithm. The core insight is that the learned neural CBF should avoid over-estimating the safety of states and actions not seen in the dataset. Our proposed method exhibits this property as can be seen in Figure 1. This objective is similar to the CQL objective, which minimizes the $Q$-values for randomly sampled actions at the dataset states in order to learn a $Q$-function that lower bounds the true one. That addresses the common problem of overestimating the $Q$-values for unseen state-action pairs in offline RL. By analogously minimizing the barrier values of states reached by random actions starting from the dataset states, CCBFs avoid overestimating the safety of unseen states.
@@
Figure 1: Learned barrier functions and trajectory rollouts.
(a) Offline dataset trajectories with an obstacle (gray) and a goal (purple) (b–d) NCBF, CCBF (ours), and iDBF, respectively. (e–j) Rollouts (blue) using three CBF-QP policies, with nominal controllers ordered left to right as PD, BC, and BC-Safe (defined in Section IV-B). For each pair of figures, the left panel shows the results of using NCBF as the CBF in CBF-QP policy and the right panel shows the results of using CCBF instead. CCBF consistently balances safety with goal-reaching by accounting for both the obstacle and the training distribution, whereas NCBF emphasizes only obstacle avoidance and iDBF is very conservative.
II Related Work
@@
Several algorithms have been proposed recently to train neural CBFs from offline datasets [10, 23, 24]. They differ in their objectives, their dataset generation process, their assumptions, and their guarantees, if any. [25] and [26] proposed learning safety filters to prevent control systems from reaching OOD states. [25] proposed an algorithm for training neural CBFs, termed in-distribution barrier functions (iDBFs), using an offline dataset of safe trajectories to prevent the control system from reaching OOD states. To train an iDBF, the algorithm first trains a BC policy using the same dataset. Then, for each state in the training dataset, the algorithm randomly samples a set of actions, and if the BC policy assigns a sampled action a probability density below a user‑defined threshold, the next state reached by that action is labeled OOD and appended to the dataset. It finally uses an existing approach for training neural CBFs (e.g., [9]) to train the iDBF using the augmented dataset, treating the InD states as safe and the OOD states as unsafe. [26] proposed Lyapunov Density Models, combining control Lyapunov functions [27] and density models [28], which are also learned from data to keep a control system from reaching OOD states. More recently, [29] proposed an algorithm for learning neural CBFs from an offline dataset consisting of a mix of labeled and unlabeled trajectories by leveraging cost-sensitive classification [30] to assign safety labels to the unlabeled demonstrations.
Several other works proposed algorithms for synthesizing CBFs from expert safe demonstrations, providing formal correctness guarantees through Lipschitz continuity and coverage-based proofs [31, 10].
In this paper, we aim to train neural CBFs from offline datasets of both safe and unsafe demonstrations that can be used as safety filters for control systems to avoid unsafe states and disincentivize reaching OOD ones while minimally affecting task progress.
@@
Safe offline RL is a branch of offline RL that aims to train policies that satisfy safety constraints while maximizing reward using offline datasets. It tackles a similar problem to ours: given a set of trajectories of a Constrained Markov Decision Process (CMDP), design an optimal and safe policy that avoids distribution shift [32]. However, it considers soft safety constraints which require the expected cost of the trajectories to remain below a user-defined threshold, in contrast with the hard constraints of avoiding failure states that we consider. Several approaches incorporated safety filters into the safe offline RL pipeline [33, 34]. A close work to ours is the recently proposed method FISOR [33], which similarly employs a hard state-wise constraint formulation for safe offline RL. It first adapts Hamilton-Jacobi reachability to learn a value function that defines the backward reachable set (BRS) of a known constraint set (aka failure set) from the offline dataset. It uses Implicit Q-learning [35] to prevent over-optimistic safety value estimation that can result from querying OOD actions in offline RL settings. It then trains a weighted BC policy that maximizes rewards within the feasible region in the state space, i.e., the complement of the backward reachable set. We instead take a different approach and assume that the states in the dataset are labeled as safe (belong to a controlled forward invariant set disjoint from the failure set) or unsafe (i.e., not safe), train a neural CBF with an additional loss term inspired from CQL [18] to avoid over-optimistic CBF values for OOD states, and test its performance as a safety filter for various nominal controllers that are learned separately from the same dataset using various safe offline RL algorithms.
III Preliminaries
In this paper, we consider nonlinear control-affine systems.
Definition 1 (Control-affine systems)
@@ III Preliminaries
A set of states $S\subseteq\mathcal{X}$ is called controlled forward invariant with respect to system (1) if there exists a control policy that makes it forward invariant, i.e., all trajectories starting in $S$ remain inside it.
We assume that there is a user-defined set of failure states, such as collisions or violations of speed limits of autonomous vehicles, that a safe controller should prevent the system from reaching. Given a controlled forward invariant set $S$ that is disjoint from the failure set, we say that the states in $S$ are safe (and $S$ is the safe set) and the states in $\bar{S}:=\mathcal{X}\SS$ are unsafe (and $\bar{S}$ is the unsafe set). Note that $\bar{S}$ contains the failure set and might also contain additional states, particularly when there are input constraints or the system has a high relative degree [16, 36]. Thus, safety is defined as the invariance to a set disjoint from the failure set. The largest safe set $S$ can be obtained using Hamilton-Jacobi (HJ) reachability analysis [36]. Unfortunately, existing algorithms for HJ reachability do not generally scale beyond few dimensions [37].
III-A Control barrier functions
Definition 2 (Control barrier functions [4])
@@ III-A Control barrier functions
A function $B:\mathcal{X}\to\mathbb{R}$ is a control barrier function for system (1) if it is continuously differentiable and satisfies: $\forall x\in\mathcal{D}\subseteq\mathcal{X}$,
$\exists u\in\mathcal{U}\text{ such that }\dot{B}(x,u)+\alpha(B(x))\geq 0,$ (2)
@@
where the super-level set $B_{\geq 0}:=\{x\ |\ B(x)\geq 0\}$ belongs to $\mathcal{D}$, and $\alpha:\mathbb{R}\to\mathbb{R}$ is a locally Lipschitz extended class-$\mathcal{K}_{\infty}$ function, i.e., it is strictly increasing and satisfies $\alpha(0)=0$.
@@
A CBF $B$ for system (1) specifies the controls that guarantee the forward invariance of $B_{\geq 0}$, as stated in the following theorem. If the set of unsafe states is disjoint from $B_{\geq 0}$ and the system starts from $B_{\geq 0}$, then the system can be kept safe by following such controls.
Theorem 1 ([4])
@@
Any Lipschitz continuous control policy $\pi:\mathcal{D}\to\mathcal{U}$ satisfying $\forall x\in\mathcal{D}$, $\pi(x)\in K_{\text{cbf}}:=\{u\in\mathcal{U}\mid\nabla B(x)(f(x)+g(x)u)+\alpha(B(x))\geq 0\}$ makes the set $B_{\geq 0}$ forward invariant.
@@
Given a reference controller $\pi_{\text{ref}}:\mathcal{X}\to\mathcal{U}$ that does not necessarily satisfy safety constraints, a safety filter using the CBF $B$ can be employed to modify its unsafe control choices. Specifically, a quadratic program (QP) can be formulated to determine the closest safe control to $\pi_{\text{ref}}(x)$ [5], resulting in what is denoted by a CBF-QP policy $\pi_{\text{safe}}$, as shown below:
$\pi_{\text{safe}}(x)$ $=\arg\min_{u\in\mathcal{U}}\lVert u-\pi_{\text{ref}}(x)\rVert^{2}$ (3)
$\quad\text{s.t. }\nabla B(x)(f(x)+g(x)u)+\alpha(B(x))\geq 0.$
Under some non-restrictive conditions, it can be proven that $\pi_{\text{safe}}$ is locally Lipschitz continuous for $x\in\text{Int}(B_{\geq 0})$, the interior of the set $B_{\geq 0}$ [5].
III-B Offline reinforcement learning
The goal of offline RL is to design optimal control policies from offline datasets without further interaction with the environment. Conservative Q learning (CQL) [18] is a prominent offline RL approach with several variants. The first variant, which is the one we are interested in here and simply call CQL for the rest of the paper, learns a $Q$ function from the offline dataset corresponding to some policy that lower bounds its actual $Q$ function pointwise for all state-action pairs, guaranteeing not over-estimating the values of state-action pairs that are not part of the dataset. Similarly, we want to learn barrier functions which do not over-estimate the safety of states not part of the dataset.
CQL modifies the $Q$-function learning iterative update for a policy $\pi$ by adding a term to minimize the $Q$-values of a set of state-action pairs, where the states are those in the dataset and the actions are sampled from some distribution $\mu(a|s)$, to the standard Bellman error term, as follows:
$\hat{Q}^{k+1}$ $\;\leftarrow\;\arg\min_{Q}\;\alpha\,\mathbb{E}_{s\sim\mathcal{D},\,a\sim\mu(a|s)}[Q(s,a)]$
$\quad+\tfrac{1}{2}\,\mathbb{E}_{s,a\sim\mathcal{D}}\Big[Q(s,a)-\hat{B}^{\pi}\hat{Q}^{k}(s,a)\Big]^{2},$ (4)
where $\alpha$ is a hyperparameter, $\mathcal{D}$ is the offline dataset, $\hat{Q}^{k+1}$ denotes the learned $Q$-function at iteration $k+1$, $\hat{Q}^{k}$ is the learned $Q$-function from the previous iteration, and $\hat{B}^{\pi}$ is the empirical Bellman operator for policy $\pi$. Theorem 3.1 in [18] establishes that the fixed point $\hat{Q}^{\pi}=\lim_{k\to\infty}\hat{Q}^{k}$ obtained by the update in equation (4) when the support of $\mu$ is a subset of the support of $\pi_{\beta}$ is a pointwise lower bound for the true $Q$-function of policy $\pi$, i.e.,
$\hat{Q}^{\pi}(s,a)\;\leq\;Q^{\pi}(s,a),\qquad\forall(s,a),$
when $\alpha$ is sufficiently large. Thus, the learned $Q$-function $\hat{Q}^{\pi}$ conservatively underestimates the true $Q$-function $Q^{\pi}$, preventing over-optimistic estimates for out-of-distribution actions. Previous research has already built a connection between control barrier functions and value functions [38, 16], so we adopt a similar analogy here to design a conservative loss for learning control barrier functions which do not over-estimate the safety of states unseen in the dataset.
III-C Problem statement
We consider the setting where an offline trajectory dataset, that is generated using a potentially unknown and unsafe reference controller, is available, along with a user-specified failure set of states. We assume that the user has domain knowledge that allows them to label some, if not all, of the states in the dataset whether they belong to a (not exactly known) controlled forward invariant set that is disjoint from the user-specified failure set, i.e., safe, or not [^1] One can only label the states which belong to the provided failure set as unsafe and leave the rest of the states in the dataset unlabeled. However, that may lead to a neural CBF with a small or empty super-level set, which represents the safe set. That limits the states at which the system can start and the resulting safety filter be effective. Alternatively, we can adapt our method to training policy neural CBFs [16], which does not require the states to be labeled as safe or unsafe to train a neural CBF with a large super-level set, but we leave that for future work.. Our objective is to train a neural CBF using the offline dataset that can be used to define the constraints in a quadratic program that acts as a safety filter that prevents the system from reaching the failure states. Consequently, it must also not over-estimate the safety of states unseen in the dataset which can result in trajectories reaching the failure set. Finally, the safety filter should not be conservative beyond what is necessary to maintain safety, i.e., as minimally altering the nominal control as possible.
IV Method